Regression equation:
P = 0.15x - 28 ( values given in thousands)
Profit = $150x - $28,000
If profit must equal $100,000, how many items need to be sold
I don't know how to put this in a formula format. I appreciate any help you can give in since algebra is new to me. Thank you so much for being there when needed. You are a godsend. Thank you! Thank you!!!!

Answer by Theo(13342)

if you are dealing in thousands, then the formula is:
P = .15x - 28
profit must be equal to 100.
set P = to 100 in the formula to get:
100 = .15x - 28
add 28 to both sides of the eqution to get:
128 = .15x
divide both sides of the equation by .15 to get:
x = 853.333333333
that's how many items need to be sold to break even.
853.33333333 * .15 - 28 = 100
since you can't sell part of an item, then you need to sell 854 items.
854 * .15 - 28 = 100.1 (you made a profit).
853 * .15 - 28 = 99.95 ( you didn't make a profit).
whether you deal in thousand or in units, you'll get the same answer.
100,000 = 150*x - 2,8000
add 28000 to both sides of the equation to get:
128,000 = 150*x
divide both sides of the equation by 150 to get:
x = 853.33333333